An artificial satellite circles the Earth in a cir
Answer #1

g = G Me / (Re + h)^2

Re + h = r

g = G Me /r^2 = 9 m/s^2

(6.67 x 10^-11 x 5.972 x 10^24)/ r^2 = 9

r = 6.653 x 10^6 m

G Me m /r^2 = m v^2 / r


v^2 / r = G Me / r^2

v^2 = 6.653 x 10^6 x 9

v = 7737.88 m/s

T = 2 pi r / v = 5402.26 sec or 1.5 hrs

2. T = 2pi r / v

v = (2 x pi x 4.22 x 10^5 x 10^3 m) / (1.77 x 24 x 3600 s)

v = 17338.25 m/s


And G M m / r^2 = m v^2 / r

G M / r = v^2

(6.67 x 10^-11) (M) = (17338.25)^2 (4.22 x 10^8)

M = 1.902 x 10^27 kg